PRECISION LASER SPECTROSCOPY OF H_{3}^{+}

\underline{HSUAN-CHEN CHEN}, Institute of Photonics Technologies, National Tsing Hua University, Hsinchu 30013, Taiwan; JIN-LONG PENG, Center for Measurement Standards, Industrial Technology Research Institute, Hsinchu 30011, Taiwan; T. AMANO, Department of Chemistry and Department of Physics and Astronomy, University of Waterloo, Waterloo, ON N2L 3G1, Canada; JOW-TSONG SHY, Institute of Photonics Technologies, National Tsing Hua University and Department of Physics, National Tsing Hua University, Hsinchu 30013, Taiwan.

The high-resolution sub-Doppler Lamb dips of the ν_{2} fundamental band transitions of H_{3}^{+} have been observed using an extended negative glow discharge tube as an ion source and a periodically poled lithium niobate optical parametric oscillator as a radiation source.^{a} The absolute frequencies of five transitions were measured to an accuracy of 250 kHz using a fiber optical frequency comb. In addition, we measured the homogeneous linewidths of these lines. Physical significance of these results will be discussed in terms of collisional processes.

aH.-C. Chen, C.-Y. Hsiao, J.-L. Peng, T. Amano, and J.-T. Shy, Phys. Rev. Lett., 109, 263002 (2012)